Solve the quadratic function by completing the square.
y=-x^2-7x-30

Answer by edjones(8007)

y=-x^2-7x-30
-x^2-7x=30
x^2+7x=-30
x^2+7x+(49/4)=(49/4)-30
x^2+7x+(49/4)=-71/4
x+7/2=+-sqrt(-71)/2
x=(-7+sqrt(71)i)/2, x=(-7-sqrt(71)i)/2
.
Ed